OpenStudy (anonymous):

I need someone to explain finding zeros of a function... I dont get it >.< Helllppp.....Solve the equation: x2 - 10x + 25 = 16.

OpenStudy (ash2326):

We have the equation \[x^2-10x+25=16\] Let's simplify this, then we'll find the zeros. Could you subtract 16 from both the sides and write what would you get? @ginlynn17

OpenStudy (anonymous):

x^2 - 10x+ 25- 16 ?

OpenStudy (anonymous):

oh wait. x^2-10x+25-16=0

OpenStudy (ash2326):

right , \[x^2-10x+25-16=0\] or \[x^2-10x+9=0\] We could solve this in two ways, using quadratic formula or by factoring. Which one do you want?

OpenStudy (anonymous):

which ones easier? :D

OpenStudy (ash2326):

Which one you are supposed to use? :P

OpenStudy (anonymous):

It doesn't matter xD ... Can you explain factoring to me in like more detail? I haven't used factoring since like 8th grade .-.

OpenStudy (ash2326):

Cool :)

OpenStudy (ash2326):

\[x^2-10x+9=0\] Standard quadratic equation is \[ax^2+bx+c=0\] Can you compare with our equation and tell what's the value of a. b and c?

OpenStudy (anonymous):

The value of a is 1, b is -10, and c is 9

OpenStudy (ash2326):

Good :) Now find the product of a and c \[a\times c\ ??\]

OpenStudy (anonymous):

It's 9

OpenStudy (ash2326):

Right :) Now you need to find the factors of 9, such that their sum is b which is -10 Can you try>?

OpenStudy (anonymous):

-9 and -1 ?

OpenStudy (ash2326):

That's great :)

OpenStudy (ash2326):

Now we write -10 x as -9x-x \[x^2-9x-x+9=0\] Now let's start taking common factors out \[x(x-9)-x+9=0\] Do you get this step?

OpenStudy (ash2326):

@ginlynn17 ???

OpenStudy (anonymous):

yeah i get the step

OpenStudy (ash2326):

Now I'll take out -1 from the other two terms \[x(x-9)-1(x-9)=0\] Let's take out (x-9) from both the terms \[(x-9)(x-1)=0\] Do you get this?

OpenStudy (anonymous):

yeah

OpenStudy (ash2326):

So, we're done either x-9=0 or x-1=0 so \[x=1, 9\] voila !

OpenStudy (anonymous):

Wait so instead of doing the work that you just did, could you just say (x-9)(x-1)=0 after you find the factors?

OpenStudy (anonymous):

And thank you so much :) You're good at explaining it. <3

OpenStudy (ash2326):

yes :)

OpenStudy (ash2326):

Welcome :)